x+19=2(x+29)
I really just don't understand the steps to take on this problem because its dealing with midsegment theorem on triangles Ex. Triangle ABC has a midsegment of DE AB=1/2DE solve for x.

Answer by stanbon(75887)

x+19=2(x+29)
------
x + 19 = 2x + 58
-----
x = -39
==============
Cheers,
Stan H.
==============